### v4

# EMAGE: Towards Unified Holistic Co-Speech Gesture Generation via Expressive Masked Audio Gesture Modeling

Abstract
We propose EMAGE, a framework to generate full-body human gestures from audio and masked gestures, encompassing facial, local body, hands, and global movements. To achieve this, we first introduce BEAT2 (BEAT-SMPLX-FLAME), a new mesh-level holistic co-speech dataset. BEAT2 combines MoShed SMPLX body with FLAME head parameters and further refines the modeling of head, neck, and finger movements, offering a community-standardized, high-quality 3D motion captured dataset. EMAGE leverages masked body gesture priors during training to boost inference performance. It involves a Masked Audio Gesture Transformer, facilitating joint training on audio-to-gesture generation and masked gesture reconstruction to effectively encode audio and body gesture hints. Encoded body hints from masked gestures are then separately employed to generate facial and body movements. Moreover, EMAGE adaptively merges speech features from the audio’s rhythm and content and utilizes four compositional VQ-VAEs to enhance the results’ fidelity and diversity. Experiments demonstrate that EMAGE generates holistic gestures with state-of-the-art performance and is flexible in accepting predefined spatial-temporal gesture inputs, generating complete, audio-synchronized results. Our code and dataset are available111*Equal Contribution. [^2] https://pantomatrix.github.io/EMAGE/
Figure 1: EMAGE. We present a Masked Audio-Conditioned Gesture Modeling framework, along with a new holistic gesture dataset, BEAT2 (BEAT-SMPLX-FLAME), for jointly generating facial expressions, local body dynamics, hand movements and global translations, conditioned on audio and a partially or completely masked gestures. The gray denotes visible gestures, and blue represents our outputs.
1 Introduction
The final stretch in Full-body Co-Speech Gesture Generation is brief yet riddled with challenges. Despite the development of various baselines, e.g., audio conditioned models for face [19, 60, 15], body [65, 25, 13], and hand movements [40, 39, 33], along with a few attempts at merged models [28, 64], the limited availability of comprehensive data and models poses an ongoing obstacle. Instead of creating results that can directly substitute for the delicate designs of animators, in this paper, we consider a framework that can incorporate partial spatial-temporal predefined gestures and autonomously complete the remaining frames in synchronization with the audio, offering advantages in applications for creating photo-realistic digital humans[63, 56, 62, 70].
To achieve this, we first require a comprehensive gesture dataset. While several datasets are available for audio-to-body [65], audio-to-body&face [28], body-to-hands [40, 39], and common actions [45, 16], integrating them poses challenges due to varying data formats and evaluation metrics across different sub-areas. This motivates us to establish a unified, community-standard benchmark for training and evaluating holistic gestures. Most close to this end, although BEAT [39] builds the most comprehensive mocap co-speech gesture dataset with diverse modalities. Despite its size and richness, the data in BEAT is not presented in a standardized format. The dataset comprises iPhone ARKit blendshape weights and Vicon skeletons, posing challenges in training to achieve a satisfied performance due to the absence of a unified representation, e.g., vertices for vertex loss [51, 20]. Additionally, conducting direct human perceptual evaluation on these diverse modalities is inherently complex. On the other hand, SMPLX [49] and FLAME [35] are common mesh standards for the academic community for other human dynamics-related tasks, e.g., action [51] and talking head generation [14]. Driven by the motivation to facilitate the sharing of knowledge across different tasks, we present BEAT-SMPLX-FLAME (BEAT2), which consists of two main components: i) Refined SMPLX body shape and pose parameters from Mosh++ with hard-coded physical priors. ii) Optimized high-quality FLAME head parameters. By integrating SMPLX’s body and FLAME’s head, BEAT2 facilitates a comprehensive training and evaluation benchmark across multiple sub-areas in the context of co-speech human animation generation.
Table 1: Comparison of Co-Speech Gesture Datasets. We summarize gesture and face datasets in talking scenarios. ‘PGT’ and ‘MC’ denote Pseudo Ground Truth and Motion Capture, respectively. The BEAT2 (Ours) dataset is the largest dataset for motion-captured data, providing holistic, academic community standard mesh-level information.
Trinity
Mocap
2017 [22]
S2G
S2G-2D
2019 [25]
Seq2Seq
TED-2D
2019 [66]
TWH
Mocap
2019 [32]
Trimodal
TED-3D
2020 [65]
VOCA
Scan
2020 [14]
MeshTalk
Scan
2021 [55]
Hahibie et al.
S2G-3D
2021 [28]
HA2G
TED-3D+
2022 [41]
BEAT
Mocap
2022 [39]
ZEEG
Mocap
2022 [24]
Yoon. et al.
TED-SMPL
2023 [44]
Talkshow
S2G-SMPL
2023 [64]
BEAT2
Ours
2024
Head - 2D - - - 3D Scan 3D Scan 3D - ARKit - - PGT-Mesh MC-Mesh
Upper Body 3D 2D 2D 3D 3D - - 3D 3D 3D 3D PGT-Mesh PGT-Mesh MC-Mesh
Hands 3D 2D - 3D - - - 3D 3D 3D 3D PGT-Mesh PGT-Mesh MC-Mesh
Lower Body 3D - - 3D - - - - - 3D 3D - - MC-Mesh
Global Motion 3D - - 3D - - - - - 3D 3D - - MC-Mesh
Duration (hours) 4 60 97 20 97 0.5 2 38 33 76 4 30 27 60
Table 2: Comparison of Co-Speech Gesture Models. We compare with previous methods for generating face or body motion trained on co-speech datasets. The first row lists their inputs, and the subsequent rows list their outputs, respectively. Different decoder designs are denoted by the initials ‘S’ for Single, ‘M’ for Multiple, and ‘C’ for Cascaded. To the best of our knowledge, EMAGE (Ours) is the first to accept audio and partially or completely masked gestures, generating full-body audio-synchronized results.
S2G
2019 [25]
TriModal
2020 [65]
B2H
2020[47]
TWH
2021[32]
Hahibie
et al.
2021[28]
DisCo
2022[36]
HA2G
2022[41]
Face
Former
2022[20]
Rhythmic
2022[4]
BEAT
CaMN
2022[39]
Diff
Gesture
2023[71]
Code
Talker
2023[60]
Talk
Show
2023[64]
DiffStlyle
Gesture
2023[61]
Body
Former
2023[48]
Lively
Speaker
2023[69]
EMAEG
2024
Input Gestures - - Body Body - - - - - - - - - - - - Masked
Head - - - - M - - S - - - S S - - - C
Upper S S - - M S C - S C S - C S S S C
Hands - - S S M S C - - C S - C S - S C
Lower + Global - - - - - - - - - - - - - S - - C
Figure 2: Comparison of Data between BEAT2 and Others. BEAT-SMPLX-FLAME presents a new mesh-level, motion-captured, holistic co-speech gesture dataset with 60h of data. Left: We compare our refined SMPLX body parameters results (denoted as Refined Mosh) with Original BEAT’s Skeleton [39], Retargeted from AutoRegPro, and initial results of Mosh++ [49]. The refined results show correct neck flexion, appropriate head and neck shape ratios, and detailed finger representations. Right: Visualization of blendshape weights from Original BEAT [39] with ARKit’s template, Wrapped based, and handcrafted optimization. Our final handcrafted FLAME blendshape-based optimization demonstrates both accurate lip movement details and natural mouth shapes.
With the holistic data from BEAT2, we aim to enhance the coherence of individual body parts while ensuring accurate cross-modal alignment between audio and motion, c.f. Figure 1. This leads to the design of Expressive Masked Audio-Conditioned Gesture Modeling, denoted as EMAGE, a spatial and temporal transformer-based framework. EMAGE first aggregates spatial features from masked body joints. Then, it adeptly reconstructs the latent space of pretrained gestures with a switchable gesture temporal self-attention and audio-gesture cross-attention. The selection of different forward paths enables modeling the gesture-to-gesture and audio-to-gesture prior separately and effectively. Once the reconstructed latent features are obtained, EMAGE decodes local face and body gestures from four compositional pretrained Vector Quantized-Variational AutoEncoders (VQ-VAEs) and decodes global translations from a pretrained Global Motion Predictor. The compositional VQ-VAEs are key to preserving audio-related motions during the reconstruction. Through these designs, EMAGE achieves state-of-the-art performance in generating body and face gestures. It is flexible to accept audio and partially initialized gestures, recovering audio-synchronized, coherent gestures. More excitingly, EMAGE showcases its further applications in joint training multiple, non-holistic gesture datasets.
Overall, our contributions are as follows:
  • We release BEAT2, a representation-unified, mesh-level dataset that leverages Moshed SMPLX-body and FLAME head parameters.
  • We propose EMAGE, a simple yet effective holistic gesture generation framework that generates coherent gestures with partial gesture and audio priors.
  • As same as previous methods, EMAGE achieves state-of-the-art performance in generating both body and face gestures with only four frames of seed gestures.
  • EMAGE showcases the use of additional non-holistic gesture datasets for training, e.g., Trinity and AMASS, to further improve the fidelity and diversity of the results, effectively leveraging data from different datasets.
2 Related Work
Co-Speech Animation Datasets are categorized into two types, c.f. Table 1: pseudo-labeled (PGT) and motion-captured (mocap). For PGT, the Speech2Gesture Dataset [25] utilizes OpenPose [11] to extract 2D poses from News and Teaching videos. Subsequent works extend the dataset to 3D pose [28] and SMPLX [64]. The TED Dataset [66], which extracts 2D pose from TED videos, has been extended to include 3D pose [65], fingers [41], and SMPLX [44] in subsequent years. Similarly, 2D, 3D landmarks, and mesh are estimated from multi-view recorded videos [59] as PGT for talking head generation. Although pseudo-labeled approaches could theoretically extract infinite data, their accuracy is limited. For instance, the average error (in mm) for the SOTA monocular 3D pose estimation algorithm [23] in the Human 3.6M dataset [30] is 33.4, while for Vicon mocap it is 0.142 [46]. For mocap datasets, Trinity [3] features one male actor with 4h of data, and TalkingWithHands [32] collects conversation scenarios for two speakers. ZEEG [24] considers one speaker with 12 styles. For the face, 3D scanning is accurate but limited in quantity due to cost, e.g., less than 3h for BIWI [21], VOCA [14], and MeshTalk [55]. This leads to a balance between performance and quantity in the ARKit dataset [50]. The aforementioned datasets address the body and face separately. On the other hand, BEAT [39] includes both 3D pose body and ARKit [8] facial blendshape weights for the first time. However, BEAT lacks mesh data.
Co-Speech Animation Models are categorized based on their outputs, c.f. Table 2. In addition to the baselines and datasets [25, 65, 14, 55, 32, 47], several frameworks for body gestures [52, 53, 2, 33] have improved the performance of the baselines. They are trained and evaluated by selecting specific upper body joints [36, 39, 4, 5, 41, 69, 48], or all body joints [71, 61]. Recent improvements in facial gestures drive the vertex by transformer and discrete face priors [20, 60]. However, these methods only address either the face or body. Applying these methods to the full body will yield sub-optimal results as audio has different correlations with face and body dynamics. Most similar to our work, Habibie et al. [28] employs a single audio encoder and multiple decoders for generating face and body gestures. Talkshow [64] demonstrates the advantage of separating the audio encoders and decoding the body and hands using quantized codes in an auto-regressive manner. However, it lacks lower body and global motion, and there is no shared information for body and facial gesture generation. Moreover, it cannot accept partially masked body hints due to the design of a fully auto-regressive model.
Masked Representation Learning was first demonstrated to be effective in Natural Language Processing, with Bert-based models [17, 31, 42] boosting the performance of learned word embeddings in downstream tasks through a combination of masked language modeling and transformer architecture. Subsequently, Masked AutoEncoders [29] expanded masked image modeling to computer vision by removing inputted image patches. This concept of masked representation learning has since been employed in other modalities, e.g., video [18, 12, 43], audio [26, 37, 38], and point cloud [68, 67, 54]. Most related to our work, MotionBERT [72] proposes a spatial-temporal transformer to learn robust motion representation for classification-based tasks by masking 2D pose. Differing from their method, we target robust motion features for conditional motion generation tasks, requiring a balance in training between multiple modalities, e.g., audio and gesture.
3 BEAT2
We introduce how to obtain unified mesh-level data, i.e., SMPLX [49] and FLAME [35] parameters, from the original BEAT [39] in this section. BEAT utilizes a Vicon motion capture system [46] and releases 78-joint skeleton-level Biovision Hierarchy (BVH) [10] motion files. Their facial capture system uses ARKit [8] with a depth camera on the iPhone 12 Pro, extracting $\mathbb{R}^{51}$ blendshape weights. These blendshapes are designed based on the Facial Action Coding System (FACS) and are widely adopted by industry novices. However, both the motion and facial data lack mesh-level details (c.f. Figure 2), e.g., shapes and vertices.
3.1 Body Parameters Initialization via Mosh++
We initialize the SMPLX body shape and pose parameters from BEAT captured markers data based on Mosh++ [49]. Given the captured markers positions $\mathbf{m}\in\mathbb{R}^{T\times K\times 3}$, predefined markers positions offset $\mathbf{d}\in\mathbb{R}^{K\times 3}$, and user-defined vertices to marker function $\mathcal{H}$, we aim to obtain body shape $\beta\in\mathbb{R}^{300}$, poses $\theta\in\mathbb{R}^{T\times 55\times 3}$, and translation parameters $\gamma\in\mathbb{R}^{T\times 3}$. The optimization uses differentiable surface vertices mapping function $\mathcal{S}(\beta,\theta,\gamma)$ and vertex normal function $\mathcal{N}(\beta,\theta_{0},\gamma_{0})$. For each frame, the latent marker $\tilde{\mathbf{m}}\in\mathbb{R}^{T\times K\times 3}$ is calculated as
$\tilde{\mathbf{m}}_{i}\equiv\mathcal{S}_{\mathcal{H}}(\beta,\theta_{i},\gamma_{i})+\mathbf{d}\mathcal{N}_{\mathcal{H}}(\beta,\theta_{i},\gamma_{i}).$ (1)
We first select 12 frames to optimize and fix $\mathbf{d}$ and $\beta$, then optimize $\theta_{i}$, $\gamma_{i}$ for $i\in(1:T)$ by minimizing the loss terms based on $\|\tilde{\mathbf{m}}_{i}-\mathbf{m}_{i}\|^{2}$, including Data Term, Surface Distance Energy, Marker Initialization Regularization, Pose and Shape Priors, Velocity Constancy, and Soft-Tissue Term (cf. supplementary materials). The overall objective function is the weighted sum of these terms to balance accuracy and plausibility.
3.2 Body Parameters Refinement
We refine body shape and pose parameters with three simple yet effective hard-coded physical rules. i) The neck and head length should approximate $1/7$ of the total length of the body. ii) Fingers, except for the thumb, should not bend backward. iii) We employ the Kolmogorov-Smirnov test, which reveals the data is similar to a normal distribution. Subsequently, we apply a Gaussian truncation approach, where all data points falling outside the $3\sigma$ range are adjusted to align with the $3\sigma$ threshold and blended with adjacent 10 frames. We compare our refined body parameters (denoted as Mosh Refined) with Original BEAT, Retargeted SMPLX, and Mosh SMPLX in Figure 2.
Figure 3: EMAGE leverages two training paths: Masked Gesture Reconstruction (MG2G) and Audio-Conditioned Gesture Generation (A2G). The MG2G path focuses on encoding robust body hints through a spatial-temporal transformer gesture encoder and cross-attention gesture decoder. In contrast, the A2G path utilizes these body hints, and separated audio encoders to decode pretrained face and body latent features. A key component in this process is a switchable cross-attention layer, crucial for merging body hints and audio features. This fusion allows the features to be effectively disentangled and utilized for gesture decoding. Once the gesture latent features are reconstructed, EMAGE employs a pretrained VQ-Decoder to decode face and local body motions. Additionally, a pretrained Global Motion Predictor is used to estimate global body translations, further enhancing the model’s capability to generate realistic and coherent gestures.
Figure 4: Details of CRA and Pretrained VQ-VAEs. Left: Content Rhythm Attention fuses speech rhythm (onset and amplitude) with content (pretrained word embeddings from text scripts) adaptively. This results in a preference for either content or rhythm in specific frames, which encourages the generation of semantical-aware gestures. Right: We pretrain four compositional VQ-VAEs by reconstructing face, upper body, hands and lower body separately to disentangle audio-agnostic gestures explicitly.
3.3 Blendshape Weights to FLAME Parameters
Given the ARKit blendshape weights $\mathbf{b}_{\text{ARKit}}\in\mathbb{R}^{T\times 51}$, we aim to derive a transformation matrix $\mathbf{W}\in\mathbb{R}^{51\times 103}$ that maps these into FLAME parameters $\mathbf{b}_{\text{FLAME}}\in\mathbb{R}^{T\times(100+3)}$, where 100 represents the number of dimensions for expression parameters and 3 for jaw movement. Due to discrepancies between the iPhone’s and FLAME’s template mesh topology, simply optimizing FLAME parameters by minimizing the geometric error between the ARKit template vertices and the wrapped FLAME vertices does not yield satisfactory results. Instead, we release a set of handcrafted blendshape templates $\mathbf{v}_{\text{template}}\in\mathbb{R}^{52}$ on FLAME following ARKit’s FACS configuration. This approach allows direct driving of the FLAME topology vertices $\mathbf{v}\in\mathbb{R}^{T\times 5023\times 3}$ from the given blendshape weights,
$\mathbf{v}=\mathbf{v}_{\text{template}}^{0}+\sum_{j=1}^{51}\mathbf{b}_{\text{ARKit},j}\cdot\mathbf{v}_{\text{template}}^{j}$ (2)
where $\mathbf{b}_{\text{ARKit},j}$ is the weight of the $j$-th ARKit blendshape and $\mathbf{v}_{\text{template}}^{j}$ is the vertex position of the FLAME model template corresponding to the $j$-th blendshape. The term $\mathbf{v}_{\text{template}}^{0}$ represents the original template vertex position of the FLAME model. We optimize $\mathbf{W}$ by minimizing $\|\tilde{\mathbf{v}}_{j}-\mathbf{v}_{j}\|_{2}$, where $\tilde{\mathbf{v}}$ is obtained from FLAME’s LBS $\mathcal{V}(\mathbf{b}_{\text{FLAME}})$. The comparisons between ARKit data, wrap-based, and our approach are shown in Figure 2.
4 EMAGE
We introduce the details of EMAGE, Expressive Masked Audio-Conditioned Gesture Modeling (c.f. Figure 3). Given gestures $\mathbf{g}\in\mathbb{R}^{T\times(55\times 6+100+4+3)}$, representing 55 joints in Rot6D, $\mathbb{R}^{100}$ FLAME parameters, $\mathbb{R}^{4}$ foot contact labels, $\mathbb{R}^{3}$ global translations, and speech audio $\mathbf{a}\in\mathbb{R}^{T\times sk}$, where $sk=sr_{\text{audio}}/fps_{\text{gestures}}$, EMAGE jointly optimizes masked gesture reconstruction and audio-conditioned gesture generation. This optimization enhances performance during inference and enables the use of partially masked gestures to complete holistic gestures. To achieve this, we first model the quantized latent space in Sec. 4.1, following [64, 60]. We then design a separated speech audio encoder using Content Rhythm Self-Attention (CRA) as detailed in Sec. 4.2. Subsequently, we learn the body hints from masked gestures via the Masked Audio Gesture Transformer, explained in Sec. 4.3. Finally, the decoding strategy for different body segments is discussed in Sec. 4.4.
4.1 Compositional Discrete Face and Body Prior
We model full-body gestures in the separated quantized latent space (c.f. Figure 4) for several reasons. Similar to [64], the body and hands improves the diversity of the results. However, we additionally separate the face and lower body due to their differing correlations with audio. i.e.,, using a single VQ-VAE to encode both upper and lower body may lead the model to overlook gestures that occur less frequently - regardless of their relation to the audio. Specifically, a single model might focus only on recovering lower body motion, neglecting elbow movements in a speaker who is constantly walking during the conversation.
The separated quantized latent space $Q=\{\mathbf{q}_{\text{f}},\mathbf{q}_{u},\mathbf{q}_{\text{h}},\mathbf{q}_{\text{l}}\}$ for face $\mathbf{g}_{\text{f}}\in\mathbb{R}^{T\times 106}$, Upper body $\mathbf{g}_{\text{u}}\in\mathbb{R}^{T\times 78}$, Hands $\mathbf{g}_{\text{h}}\in\mathbb{R}^{T\times 180}$, and Lower body $\mathbf{g}_{\text{l}}\in\mathbb{R}^{T\times(54+4)}$ are from four VQ-VAEs. The VQ-VAE for each is optimized by jointly optimizing the following loss terms,
$q_{i}=\operatorname*{arg\,min}_{q_{i}\in\mathbf{q}}\|z_{j}-q_{i}\|^{2}$ (3)
$\mathcal{L}_{\text{VQ-VAE}}=$ $\mathcal{L}_{rec}(\mathbf{g},\hat{\mathbf{g}})+\mathcal{L}_{\text{vel}}(\mathbf{g}^{\prime},\hat{\mathbf{g}}^{\prime})+\mathcal{L}_{\text{acc}}(\mathbf{g}^{\prime\prime},\hat{\mathbf{g}}^{\prime\prime})$
$+\|\text{sg}[\mathbf{z}]-\mathbf{q}\|^{2}+\|\mathbf{z}-\text{sg}[\mathbf{q}]\|^{2},$ (4)
where $\mathbf{z}$ is the encoded $\mathbf{g}$ with a temporal window size $w=1$. $\mathcal{L}_{\text{rec}}$, $\mathcal{L}_{\text{vel}}$, and $\mathcal{L}_{\text{acc}}$ are Geodesic [57] and L1 loss, sg is a stop gradient operation, we set the weight of the commitment loss [58] (the last term) to 1 in this paper.
4.2 Content and Rhythm Self-Attention
Given the speech audio $\mathbf{s}$, inspired by [5], we employ onset $\mathbf{o}$ and amplitude $\mathbf{a}$ as explicit audio rhythm, alongside the pretrained embeddings [9] $\mathbf{e}$ from transcripts as content. Different from previous approaches, which typically add the rhythm and content features, we leverage self-attention to merge these features adaptively. This approach is driven by the observation that for specific frames, the gestures are more related to content (semantic-aware) or rhythm (beat-aware). The rhythm and content features are first encoded into time-aligned features $\mathbf{r}_{1:T}$ and $\mathbf{c}_{1:T}$, using a Temporal Convolutional Network (TCN) and linear mapping, respectively. For each time step $t\in\{1,\dots,T\}$, we merge the rhythm and content features by:
$\mathbf{f}_{1:T}$ $=\alpha\times\mathbf{r}_{1:T}+(1-\alpha)\times\mathbf{c}_{1:T},$ (5)
$\alpha$ $=\text{Softmax}(\mathcal{AT}(\mathbf{r}_{1:T},\mathbf{c}_{1:T})),$
where $\mathcal{AT}$ is a 2-layer MLP. We apply two separate CRA encoders for the face and body.
Figure 5: Comparison of Forward Path Designs. Straightforward fusion module (a) merges audio features without refined body features and recombines audio features based only on position embedding. The Self-Attention decoder module (b), adopted in previous MLM models [17, 31], is limited for tasks requiring auto-regressive inference. Our design (c) considers effective audio feature fusion and auto-regressive decoding.
4.3 Masked Audio-Conditioned Gesture Modeling
We propose a Masked Audio Gesture Transformer to leverage different training paths (cf. Fig. 5 for the motivation behind the architecture designs). Given temporal and spatially masked gestures $\overline{\mathbf{g}}\in\mathbb{R}^{T\times 337}$, we first replace the masked tokens with learned masked embeddings $e_{\text{mask}}\in\mathbb{R}^{256}$, as the value zero still represents specific motion content, e.g., a T-pose. We linearly increase the ratio of masked joints and frames from 0 to 95% according to the training epochs.
A spatial convolutional $\mathcal{SC}$ encoder initially summarizes the spatial information, and compress the spatial feature into $\mathbb{R}^{T\times 256}$. We then employ a temporal self-attention (without feed-forward) $\mathcal{TSA}$ to refine the summarized spatial features. A straightforward Teacher-Force temporal cross-attention Transformer decoder $\mathcal{TCAT}$ is adopted for gesture reconstruction. It takes in the $\overline{\mathbf{g}}$ and $e_{\text{mask}}$, outputting the reconstructed latent $\hat{\mathbf{l}}_{\text{mg2g}}$,
$\mathbf{h}=\mathcal{TSA}(\mathcal{SC}(\overline{\mathbf{g}}+\mathbf{p}_{s})+\mathbf{p}_{t}),$ (6)
where $\mathbf{h}\in\mathbb{R}^{T\times 512}$ represents encoded body hints, $\mathbf{p}_{s}$ is a learned spatial position embedding, and $\mathbf{p}_{t}$ is the sum of a learned speaker embedding and PPE [20].
$\hat{\mathbf{q}}_{\text{mg2g}}=\mathcal{TCAT}(\mathbf{h}+\mathbf{p}_{t}).$ (7)
Then, we minimize the L1 distance in the latent space,
$\mathcal{L}_{\text{mg2g}}=\|\hat{\mathbf{q}}_{\text{mg2g}}-\mathbf{q}\|.$ (8)
Masked gesture reconstruction encodes effective body hints, with the key being to leverage these body hints for gesture generation. We employ a selective fusion for audio and body hints via a temporal cross-attention $\mathcal{TCA}$. Subsequently, we use the merged audio-gesture features for audio-conditioned gesture latent reconstruction,
$\hat{\mathbf{q}}_{\text{a2g}}=\mathcal{TCAT}(\mathcal{TCA}(\mathbf{h}+\mathbf{p}_{t},\mathbf{f}_{\text{body}}),\overline{\mathbf{g}}+\mathbf{p}_{t}).$ (9)
We optimize both the latent code class classification cross-entropy loss $\mathcal{L}_{\text{a2g-rec}}$ and the latent reconstruction loss $\mathcal{L}_{\text{a2g-cls}}$ to encourage the results’ diversity.
4.4 Face and Translation Decoding
Considering that the face is weakly related to body motion, applying the same operation – recombining the audio features based on body hints – is not reasonable. Therefore, we directly concatenate the masked body hints with the audio features for the final decoding of facial latent,
$\hat{\mathbf{q}}_{\text{f}}=\mathcal{TCAT}(\mathbf{f}_{\text{face}}\oplus\mathbf{h},\mathbf{p}_{t}).$ (10)
Once we have obtained the local lower body motion $\tilde{\mathbf{g}}_{l}\in\mathbb{R}^{T\times(54+4)}$ from the VQ-Decoder, we estimate the global translations $\tilde{\mathbf{t}}\in\mathbb{R}^{T\times 3}$ with a pretrained Global Motion Predictor $\tilde{\mathbf{t}}=\mathcal{G}(\tilde{\mathbf{g}}_{l})$, which significantly reduces foot sliding.
Table 3: User Preference Win Rate between Existing Datasets. ‘Top.’ denotes the topology of the mesh. The results show our BEAT2 dataset outperforms the existing PGT dataset [64] in both body and face aspects. It also performs slightly better on body and lower on face when compared with previous mocap [45] and 3D scan-based datasets [14], respectively.
Top.-B Top.-F Body Face
VOCA[14] - FLAME - 38.3 ± 5.63
AMASS[45] SMPLX FLAME 42.0 ± 3.60 -
TalkShow[64] SMPLX SMPLX 14.4 ± 2.19 26.1 ± 6.42
BEAT2 (Ours) SMPLX FLAME 43.6 ± 3.38 35.7 ± 5.91
Figure 6: Subjective Examples. Top to down: GroundTruth, Generated W/O Body Hints, Generated With Body Hints, Visible Body Hints, EMAGE generates diverse, semantic-aware and audio synchronized body gestures, e.g. raise the both hands for “spare time”, relaxing gestures for “hike in nature”. Besides, as shown in the third and bottom rows, EMAGE is flexible to accept non-audio-synchronized body hints on any frames and joints, to guide or customize the generated gestures explicitly, e.g., repeat a similar motion for raising hands, change the walking orientations, etc. Note that in this figure, The gray joints of generated results are not copies of visible hints.
5 Experiments
We separate the evaluation into two categories: dataset quality and model ability. After removing data with low finger quality from BEAT, BEAT2 is reduced to 60 hours. We further split it into BEAT2-Standard (27 hours) and BEAT2-Additional (33 hours), based on the type of speech and conversation sections [39]. While acted gestures in speech sections are more diverse and expressive, spontaneous gestures in conversational sections tend to be more natural yet less varied. We report the results on BEAT2-Standard Speaker-2 with an 85%/7.5%/7.5% train/val/test split.
5.1 Dataset Quality Evaluation
We compare our dataset with the state-of-the-art pseudo ground truth (PGT) dataset, Talkshow [64], for both face and body. Additionally, we compare it with the AMASS [45] dataset for the body and VOCA [14] for the face as references. Due to the varying lengths of sequences in each dataset, we sample 100 comparison pairs, each with an equal duration ranging from 2 to 4 seconds. Each user evaluates a random set of 40 pairs during a 10-minute session by selecting the sequence they consider to have the best captured quality. In total, 60 users were invited to participate. It is important to note that users are instructed to compare only the upper body results, as Talkshow contains only the upper body. The results are shown in Table 3.
5.2 Model Ability Evaluation
We report Table 4 on BEATv2 and ablation studies on BEATv1.3, an earlier release version of BEATv2. Compared to BEATv1.3, BEATv2 refines the facial data by animators and filters the hand data by human annotators.
Table 4: Quantitative evaluation on BEATv2. We report FGD ×$\times 10^{-1}$, BC ×$\times 10^{-1}$, Diversity, MSE ×$\times 10^{-8}$, and LVD ×$\times 10^{-5}$. For body gestures, EMAGE significantly improves the FGD, indicating that the generated results are closer to the GT. This shows the effectiveness of body hints from masked gesture modeling.
FGD $\downarrow$ BC $\uparrow$ Diversity $\uparrow$ MSE  $\downarrow$ LVD  $\downarrow$
FaceFormer[20] - - - 7.787 7.593
CodeTalker[60] - - - 8.026 7.766
S2G[25] 28.15 4.683 5.971 - -
Trimodal[65] 12.41 5.933 7.724 - -
HA2G[41] 12.32 6.779 8.626 - -
DisCo[36] 9.417 6.439 9.912 - -
CaMN[39] 6.644 6.769 10.86 - -
DiffStyleGesture[61] 8.811 7.241 11.49 - -
Habibie et al.[28] 9.040 7.716 8.213 8.614 8.043
TalkShow[64] 6.209 6.947 13.47 7.791 7.771
EMAGE (Ours) 5.512 7.724 13.06 7.680 7.556
Table 5: User Preference Win Rate On Generated Results. The results indicate that our generated outcomes are perceived as more realistic and believable, with a 14% and 23% higher user preference for body and face gestures, respectively.
Holistic Body Face
Habibie et al.[28] 12.4 ± 3.70 15.9 ± 6.49 10.8 ± 3.19
TalkShow[64] 34.9 ± 5.79 40.4 ± 8.22 33.2 ± 6.03
EMAGE (Ours) 52.7 ± 7.91 44.7 ± 8.68 56.0 ± 7.80
Metrics
We adopt FGD[65] to evaluate the realism of the body gestures. Then, we measure Diversity[33] by calculating the average L1 distance between multiple body gesture clips, and use BC [34] to assess speech-motion synchrony. For the face, we calculate the vertex MSE[60] to measure the positional distance and the vertex L1 difference LVD [64] between the GT and generated facial vertices.
Compared Methods
We first compare our method with representative state-of-the-art methods in body gesture generation [25, 65, 41, 36, 39, 61] and talking head generation [20, 60] by reproducing their methods for body and face, respectively. In addition, we reproduce two previous state-of-the-art holistic pipelines, Habibie et al. [28] and Talkshow [64], whose original implementations are limited to the upper body. We adds a lower-body decoder to Habibie’s method and a lower-body VQ-VAE to Talkshow’s method.
5.2.1 Quantitative and Qualitative Analysis
As shown in Table 4, with a four-frame seed pose, our method outperforms previous SOTA algorithms. For qualitative results, we first present subjective evaluations in Figures 6 and G. Furthermore, we conduct a perceptual user study. Maintaining the same setup of 60 users, each user evaluates 40 pairs of 10-second results. Users decide which result is more believable as the win rate, with results shown in Table 5.
Figure 7: Comparison of Generated Facial Movements. Compared with previous state-of-the-art talking face generation methods faceformer[20] and codetalker[60], holistic gestures generation methods habibie et al.[28] and Talkshow [64], notes that in BEAT2, codetalker higher MSE (lower is better) but better subjective results as discrete face prior. but due to our method present more accurate lip motions by leveraging the advantage of discrete face prior and masked body hints.
5.2.2 Ablation Analysis
Performance of Baseline
As shown in Table 6, we start with a teacher-force transformer-based baseline. This baseline, inspired by FaceFormer [20], adopts a 1-layer cross-attention transformer decoder and replaces the audio features from Wav2Vec2 [6] with our custom TCN [7] and trainable word embeddings [9].
Effect of Multiple VQ-VAEs
Simply applying one VQ-VAE [58, 27] for full-body movements, including the face, decreases performance in facial movements. This is because the VQ-VAE is trained to minimize the average loss for the full body, and some speakers’ most frequent movements are not related to the audio. Implementing separated VQ-VAEs allows the model to better leverage the advantages of discrete priors.
Effect of Content Rhythm Self-Attention
Adaptive fusion of rhythm and content features shows improvement in both FGD and Alignment. It selectively focuses the current motion more on rhythm or content features based on the training data’s distribution. Moreover, we observed more semantic-aware results when applying CRA.
Effect of Masked Body Gesture Hints
The improvement across all objective metrics demonstrates that our model effectively leverages spatial-temporal gesture priors, reducing the likelihood of incorrect gesture sampling. Furthermore, masked gesture modeling is key to enabling the network to accept customized inputs in the middle frames.
Table 6: Abliation Analysis on BEATv1.3
FGD $\downarrow$ BC $\uparrow$ Diversity $\uparrow$ MSE $\downarrow$ LVD $\downarrow$
Ground Truth 0 6.896 13.074 0 0
Reconstruction 3.913 6.758 13.145 0.841 6.389
Baseline 13.080 6.941 8.3145 1.442 9.317
+ VQVAE 9.787 6.673 10.624 1.619 9.473
+ 4 VQVAE 7.397 6.698 12.544 1.243 8.938
+ CRA 6.833 6.783 12.676 1.186 8.715
+ Masked Hints 5.423 6.794 13.057 1.180 9.015
Table 7: Training EMAGE on Multiple Datasets. EMAGE demonstrates flexibility by training on multiple datasets, even when only a subset of holistic gestures is available. This approach further improves the objective metrics on the BEATv1.3 test set.
body hands face FGD $\downarrow$ BC $\uparrow$ Diversity $\uparrow$
BESTX 5.423 6.794 13.057
+ Trinity[22] 5.319 6.843 13.346
+ AMASS[45] 5.174 6.769 14.318
5.2.3 Ability of Multi-Dataset Training.
We demonstrate that EMAGE can effectively combine multiple non-holistic datasets for training by jointly training with the retargeted Trinity [22] and AMASS [45], accepting only the upper body and audio pairs from Trinity and only body and hands from AMASS. Separate VQ-VAEs were trained for BEAT2, Trinity, and AMASS, with separate MLP heads implemented for codebook classification. The results in Table 7 show that incorporating data improves performance on the BEATv1.3 test set.
6 Conclusion
In this work, we present EMAGE, a framework designed to accept partial gestures as input for completing audio-synchronized holistic gestures. It demonstrates that leveraging masked gesture reconstruction can significantly enhance the performance of audio-conditioned gesture generation. Furthermore, the design of EMAGE enables the training on multiple datasets, further improving performance. Alongside EMAGE, we release BEAT2, the largest mesh dataset adhering to academic standards in both face and body formats. We hope that BEAT2 will contribute to knowledge and model sharing across various subareas.
7 Conflict of Interest Disclosure
While You Zhou, Xuefei Zhe, Naoya Iwamoto, and Bo Zheng are employees of Huawei Tokyo Research Center, this work was done on their own time with the approval of their employer.
MJB has received research gift funds from Adobe, Intel, Nvidia, Meta/Facebook, and Amazon. MJB has financial interests in Amazon and Meshcapade GmbH. While MJB is a co-founder and Chief Scientist at Meshcapade, his research in this project was performed solely at, and funded solely by, the Max Planck Society.
Supplementary Material
This supplemental document contains seven sections:
  • Evlauation Metrics (Section A).
  • BEAT2 Dataset Details (Section B).
  • Baselines Reproduction Details (Section C).
  • Settings of EMAGE (Section D).
  • Visualization Blender Addon (Section E).
  • Training time (Section F).
  • Importance of lower body motion (Section G).
Appendix A Evaluation Metrics
Fréchet Gesture Distance
A lower FGD, as referenced by [65], indicates that the distribution between the ground truth and generated body gestures is closer. Similar to the perceptual loss used in image generation tasks, FGD is calculated based on latent features extracted by a pretrained network,
$\operatorname{FGD}(\mathbf{g},\hat{\mathbf{g}})=\left\|\mu_{r}-\mu_{g}\right\|^{2}+\operatorname{Tr}\left(\Sigma_{r}+\Sigma_{g}-2\left(\Sigma_{r}\Sigma_{g}\right)^{1/2}\right),$
where $\mu_{r}$ and $\Sigma_{r}$ represent the first and second moments of the latent features distribution $z_{r}$ of real human gestures $\mathbf{g}$, and $\mu_{g}$ and $\Sigma_{g}$ represent the first and second moment of the latent features distribution $z_{g}$ of generated gestures $\hat{\mathbf{g}}$. We use a Skeleton CNN (SKCNN) based encoder [1] and a Full CNN-based decoder as the autoencoder pretrained network. The network is pretrained on both BEAT2-Standard and BEAT2-Additional Data. The choice of SKCNN over a Full CNN encoder is due to its enhanced capability in capturing gesture features, as indicated by a lower reconstruction MSE loss (0.095 compared to 0.103).
L1 Diversity
A higher Diversity [33] indicates a larger variance in the given gesture clips. We calculate the average L1 distance from different $N$ motion clips as follows:
[…]
where $p_{t}$ represents the position of joints in frame $t$. We calculate diversity on the entire test dataset. Additionally, to compute joint positions, the translation is set to zero, implying that L1 Diversity is focused solely on local motion.
Beat Constancy (BC)
A higher BC, as defined by [34], suggests a closer alignment between the gesture’s rhythm and the beat of the audio. We identify the beginning of speech as the audio beat and consider the local minima of the velocity of the upper body joints (excluding fingers) as the motion beat. The synchronization between audio and gesture is computed in the following manner:
$\text{BC}=\frac{1}{g}\sum_{b_{g}\in g}\exp\left(-\frac{\min_{b_{a}\in a}\left\|b_{g}-b_{a}\right\|^{2}}{2\sigma^{2}}\right),$
where $g$ and $a$ represent the sets of gesture beat and audio beat, respectively.
Appendix B BEAT2 Dataset Details
Statistics
The original BEAT dataset, as described by [39], contains 76 hours of data for 30 speakers. We exclude speakers 8, 14, 19, 23, and 29, which account for 16 hours of data, due to noise in the finger data, leaving 60 hours of data for 25 speakers (12 female and 13 male). The speech and conversation portions are categorized into BEAT2-standard and BEAT2-additional, containing 27 and 33 hours respectively. We adopt an 85%, 7.5%, and 7.5% split for BEAT2-standard, maintaining the same ratio for each speaker. BEAT2-additional is utilized for further improving the network’s robustness. The results presented in this paper are based on training with BEAT2-standard speaker-2 only. The dataset includes 1762 sequences with an average length of 65.66 seconds per sequence. Each recording in a sequence is a continuous answer to a daily question. Additionally, we report a comparison between TalkShow [64] and our dataset in terms of Diversity and Beat Constancy (BC), as shown in Table 8.
Table 8: Diversity and BC Comparisons. The local and global diversity refers to the variance in joint positions with and without global translations, respectively.
BC $\uparrow$ Diversity-L $\uparrow$ Diversity-G $\uparrow$
TalkShow [64] 6.104 5.273 5.273
BEAT2 (Ours) 6.896 13.074 27.541
Loss Terms of MoSh++
MoSh’s optimization involves loss functions including a Data Term, Surface Distance Energy, Marker Initialization Regularization, Pose and Shape Priors, and a Velocity Constancy Term, which are described as follows:
  • Data Term ($E_{D}$): Minimizing the squared distance between simulated and observed markers. In the given context, the $\tilde{M},\beta,\Theta$, and $\Gamma$ represent the latent markers, body shape, poses, and body location respectively:
    $E_{D}(\tilde{M},\beta,\Theta,\Gamma)=\sum_{i,t}||\hat{m}(\tilde{m}_{i},\beta,\theta_{t},\gamma_{t})-m_{i,t}||^{2}.$ (14)
  • Surface Distance Energy ($E_{S}$): Ensuring markers maintain prescribed distances from the body surface:
    $E_{S}(\beta,\tilde{M})=\sum_{i}||r(\tilde{m}_{i},S(\beta,\theta_{0},\gamma_{0}))-d_{i}||^{2}.$ (15)
  • Marker Initialization Regularization ($E_{I}$): Penalizing deviations of estimated markers from initial positions:
    $E_{I}(\beta,\tilde{M})=\sum_{i}||\tilde{m}_{i}-v_{i}(\beta)||^{2}.$ (16)
  • Pose and Shape Priors: Penalizing deviations from mean shape and pose:
    $E_{\beta}(\beta)=(\beta-\mu_{\beta})^{T}\Sigma^{-1}_{\beta}(\beta-\mu_{\beta}),$ (17)
    $E_{\theta}(\Theta)=\sum_{t}(\theta_{t}-\mu_{\theta})^{T}\Sigma^{-1}_{\theta}(\theta_{t}-\mu_{\theta}).$ (18)
  • Velocity Constancy Term ($E_{u}$): Reducing marker noise and ensuring movement consistency:
    $E_{u}(\Theta)=\sum_{t=2}^{n}||\theta_{t}-2\theta_{t-1}+\theta_{t-2}||^{2}.$ (19)
The overall objective function is the weighted sum of these terms, balancing accuracy and plausibility:
$E(\tilde{M},\beta,\Theta,\Gamma)=\sum_{\omega\in\{D,S,\theta,\beta,I,u\}}\lambda_{\omega}E_{\omega}(\cdot).$ (20)
More details and pseudo code of the head and neck shape optimization are available in the code release.
Details of FLAME Parameter Optimization
To animate a face using the SMPL-X model with ARKit parameters from the BEAT dataset, we estimate FLAME expression parameters by minimizing the geometric error between an animated ARKit and FLAME model. Addressing the optimization challenges posed by differing mesh structures, we construct an ARKit-compatible FLAME model utilizing Faceit, a Blender add-on tailored for crafting ARKit blendshapes. By driving the ARKit-aligned FLAME model with each set of ARKit parameters from the BEAT dataset, we obtain original FLAME expression parameters by minimizing the L2 distance loss between equivalent vertices. Finally, the optimized FLAME expression parameters can be directly applied to SMPL-X. For facial identity parameters, we preserve the same identity parameters on SMPL-X after body fitting with MoSh++ [45].
Appendix C Baselines Reproduction Details
Number of Joints
All baseline methods output full-body joint rotations represented by $\mathbf{g}\in\mathbb{R}^{T\times(55\times 6)}$ and, in addition to rotations, they decode global translations $\in\mathbb{R}^{T\times 3}$. To provide a thorough comparison, we present subjective results for both the upper body (excluding global motion) and the full body.
Autoregressive Training
We observe that autoregressive training/inference-based models, such as FaceFormer and CodeTalker [20, 60], perform worse than non-autoregressive methods. In non-autoregressive settings, only positional embedding is used as input for cross-attention to audio features, particularly when training with Rot6D and axis-angle representations. The network architecture of FaceFormer and CodeTalker is based on transformers and was initially proposed for training with the representation of vertex offsets. As shown in Table 9, we find that non-autoregressive training improves performance with FLAME’s parameters. The results in this paper are obtained using a non-autoregressive training approach. Non-autoregressive training techniques have also been employed in the training of EMAGE.
Table 9: Vertex Errors (MSE) with Different Training Methods. ‘FF’ and ‘CT’ refer to FaceFormer [20] and CodeTalker [60], respectively. ‘TF’, ‘AR’, and ‘NonAR’ represent Teacher-Force, AutoRegressive, and Non-AutoRegressive training, respectively. We train on the VOCA dataset with a vertex loss, and BEAT2 with a FLAME parameter loss combined with a vertex loss. Results indicate that the same method performs differently when using the two representations; in BEAT2, non-autoregressive training demonstrates superior performance. The average MSE is calculated on 5023 and 10475 vertices for VOCA and BEAT2, respectively:
FF-TF FF-AR FF-NonAR CT-TF CT-AR CT-NonAR
VOCA (x10-7) 6.636 6.023 6.138 7.914 7.637 7.541
BEAT2 (x10-7) 2.167 3.704 1.195 2.079 4.120 1.243
Adversarial Training
We omit the adversarial training in Speech2Gesture [25], CaMN [39], and Habibie et al [28], because their outputs with adversarial training show noticeable jitter, even when we increase the weight for the velocity loss. Similar effects are also observed in training with 3D data for Speech2Gesture [25], as reported in the study by [33].
Lower Body VQ-VAE for TalkShow
We introduce an additional VQ-VAE for TalkShow, utilizing their autoregressive (AR) model to jointly predict the class index of the upper body, hands, and lower body. The global translations are encoded in conjunction with lower body joints.
Appendix D Settings of EMAGE
Training
[…]
Structure of VQ-VAE
[…]
Global Motion Predictor
We train the Global Motion Predictor using an architecture that mirrors the CNN-based structure of our VQ-VAE’s encoder and decoder The input consists of local motions and predicted foot contact labels $\in\mathbb{R}^{T\times 334}$, and it outputs global translations $\in\mathbb{R}^{T\times 3}$.
Appendix E Visualization Blender Add-on
For straightforward visualization of our BEAT2 dataset, we utilize the SMPL-X Blender add-on [49]. As the latest SMPL-X add-on does not support the full range of facial expressions for SMPL-X, we extract 300 expression meshes from the original SMPL-X model and added them as individual blendshape targets into the SMPL-X model within the Blender add-on.
Appendix F Training time
We report the training time on a single L4, V100 and 4090 with a batch size (BS) of 64 for the best performance.
1-speaker 25-speaker
EMAGE 1-epoch 400-epoch 1-epoch 100-epoch Mem. BS
L4 (24G) 239s 26.5h 3197s 89.6h 20.1G 64
V100 (32G) 155s 17.2h 2073s 58.1h 20.1G 64
4090 (24G) 72s 8.0h 963s 27.1h 20.1G 64
Addtionally, pretraining of the 5 $\times$ VQVAEs for face, hands, upper body, lower body, and global motion would take 22.4 hours on 5 $\times$ 4090 GPUs.
1-speaker 25-speaker
VQVAEs $\times$ 1 1-epoch 700-epoch 1-epoch 100-epoch Mem. BS
L4 (24G) 200s 39.5h 2760s 74.4h 13.8G 64
V100 (32G) 131s 25.5h 1727s 48.0h 13.8G 64
4090 (24G) 61s 11.9h 802s 22.4h 13.8G 64
Appendix G Importance of lower body motion
Lower body motion allows gestures semantically aligned with the content of the audio to achieve more vivid and impressive results, e.g., “hiking in nature” with a walking gesture, “playing football” with a kicking motion; see figure below. Compared with the upper body, it is more weakly related to the audio, but it still has connections in the above cases.
In the implementation of EMAGE, we first obtain the latents of different body components with separate MLPs. Then, the lower body motion decoder leverages all latents of “audio”, “upper body”, and “hands” for cross-attention based lower-body motion decoding. We have also observed that decoding directly from audio increases diversity but reduces the coherence of the results on BEATv1.3.
[…]
audio only 6.209 6.683 13.714 1.183 8.788
audio + upper + hands 5.423 6.794 13.075 1.180 8.715